OpenStudy (anonymous):

x^2-x+3=0 solve for x

OpenStudy (saifoo.khan):

ishan singh is here!

OpenStudy (anonymous):

\[1/2(1+i \sqrt{11}),1/2(1-1\sqrt{11})\]

OpenStudy (anonymous):

I believe you'd just have to do the quadratic formula for that one.

OpenStudy (anonymous):

that was my answer

OpenStudy (anonymous):

second part should be i not 1 after the minus

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac} }{2a}\] when you do that you get imaginary roots

OpenStudy (anonymous):

Hi Saifoo

OpenStudy (anonymous):

is my answer wrong?

OpenStudy (anonymous):

You are correct.

OpenStudy (anonymous):

no...your is right

OpenStudy (saifoo.khan):

hello eshan.

OpenStudy (anonymous):

cool medals all around lol

OpenStudy (saifoo.khan):

HEHE.